Based on a predicted level of production and sales of 15,000 units, a company anticipates reporting operating income of $23,000 after deducting variable costs of $90,000 and fixed costs of $7,000.Based on this information, the budgeted amounts of fixed and variable costs for 18,000 units would be:

$8,400 of fixed costs and $108,000 of variable costs.

$7,000 of fixed costs and $108,000 of variable costs.

$7,000 of fixed costs and $102,000 of variable costs.

$7,000 of fixed costs and $90,000 of variable costs.

$8,400 of fixed costs and $90,000 of variable costs.
Answer #1

Variable cost per unit=90,000/15000=$6 per unit

Hence total variable cost for 18000 units

=($6*18000)=$108,000

Total fixed cost for 18000 units=$7000

Hence the correct option is:

$7,000 of fixed costs and $108,000 of variable costs.

NOTE:Total fixed costs and variable cost per unit do not change with change in units